# See pages - of ArxivIEEE_CopyrightPage_Submission.pdf

### v1

@@ Abstract
In this work, the effect of the measurement position on measurements of H(3) of a new active eye lens dosemeter prototype based on the Dosepix detector is examined. A comparison between measuring directly in front of the eye and measuring at the side of the head of an Alderson phantom showed no significant influence on the resulting H(3) for different radiation qualities and angles. In addition, to account for the absorption effect of radiation safety glasses, pieces of lead glass were attached to the front and side of the dosemeter. The challenging geometry of the human face, when additionally equipped with absorbing glasses and irradiated at different angles of radiation incidence, led to a complex irradiation setup, whereby the dosemeter cannot be placed at the real point of interest, i.e. directly in front of the eye. Corresponding effects and consequences for radiation protection measurements have been investigated by using a human like Alderson head phantom as well as thermoluminescent dosemeters (TLDs) and the active eye lens dosemeter prototype based on the Dosepix detector. For specific angles, the radiation was able to bypass the radiation safety glasses and lead glass pieces, leading to an increase in the measured H(3). Compared to TLDs behind radiation safety glasses, measurements with the lead glass shielded prototype resulted in lower values for H(3). Furthermore, the results did not reproduce previous findings where larger dose values were found for the Dosepix behind lead glass pieces than for the TLDs behind radiation safety glasses. A possible reason might be that the dimensions of the lead glass pieces are not representative of the radiation safety glasses in front of the eye but, ultimately, it is not yet clear what the main reason for the deviation is. Therefore, it is advisable to test the same methodology in future investigations with other eyewear models and lead glass pieces to investigate whether similar behaviors occur.
@@
The most important quantity in the context of eye lens dosimetry is the operational quantity . The International Commission on Radiation Units (ICRU) defines it as the equivalent personal dose at a depth of 3 mm in ICRU soft tissue [6, 7]. This quantity is used to estimate the dose deposited in the lens of the eye [8], i.e. the equivalent dose in the lens of the eye.
A prototype of an active dosemeter for the eye lens was developed at the Erlangen Centre for Astroparticle Physics (ECAP) as part of Friedrich-Alexander-Universität Erlangen-Nürnberg (FAU) [9]. This dosemeter is based on the hybrid pixel detector Dosepix and provides excellent results in measurements. However, it is unclear how the behavior of the dosemeter changes under more realistic conditions. So far, the eye lens dosemeter prototype has only been tested at the center front of a water cylinder phantom [9] in accordance with corresponding type test standards by IEC such as IEC 61526:2024 [10]. However, this does not represent the routine use since the dosemeter will be worn at the side of the head. This work examines the differences between those two measurement positions. Moreover, medical staff often use radiation safety glasses, which were not taken into account in previous investigations. In order to reproduce the shielding effect of such glasses, this work investigates the effectiveness of using lead glass pieces attached to the dosemeter.
@@ III The Eye Lens dosemeter
The prototype of the active eye lens dosemeter used for the measurements contains a Dosepix detector placed 2.5 mm behind the front wall of the 3D-printed casing. To ensure a stable position and prevent possible damage to the sensor, a PMMA ring with a diameter of 18 mm and a thickness of 2.5 mm is fixed around the detector. The read-out hardware is connected to the Dosepix detector via a flat cable. Detailed information as well as the results of performance tests are provided in [9].
During the measurement process, each detector pixel compares incoming energy depositions to 16 freely adjustable energy bins and sorts them accordingly. The lowest bin edge is set to 12 keV. The highest bin is used as an overflow bin and contains all events for which the deposited energy exceeds 150 keV. The counts in each bin are summed up to the total number of entries , which is then multiplied with a corresponding conversion factor . The weighted sum over all 16 bins determines the equivalent dose of the eye lens .
The conversion factors were determined by a combination of measurements and radiation transport simulations and yield reliable results for photon radiation fields with energies up to 250 keV [9]. The statistical uncertainty of the dose is estimated from Poisson statistics with the law of uncertainty propagation under the assumption of independent [15, 16].
@@ IV-A Methods
Fig. 1 shows an image of the setup for the measurements at the facilities of Physikalisch-Technische Bundesanstalt (PTB). To compare the values of measured on the eye and at the side of the head, two Dosepix detectors (numbered as Dpx 1 and Dpx 2 respectively) were placed on an Alderson phantom [19]. This phantom allows for a more representative position of the dosemeter as it would be worn by the user. It should be noted that neither of the two detectors is installed in a dosemeter prototype, but is directly connected to the read-out hardware. To minimize the amount of material in front of the phantom, a connector extension is installed between Dosepix and its read-out hardware. Both detectors are in the same plane perpendicular to the beam axis. The reference point according to ISO 29661 [20] is set to the nasal root of the Alderson phantom at 2.5 m distance to the X-ray tube. The detector at the side is placed at a distance of 7.0 cm to the reference point, the detector on the eye at a distance of 2.5 cm.
The setup was irradiated with radiation qualities of the N-series from N-20 to N-300 according to ISO 4037-1 [21]. This series of radiation qualities was chosen for comparability reasons with previous measurements [9, 16]. To account for possible differences between the two detectors, the measurements were repeated where the detectors, including the respective hardware, were exchanged. Additionally, the angle of irradiation was varied between 0∘ to , and for the measurements with N-30, N-60, N-100 and N-200. Negative angles denote a clockwise rotation and positive angles a counterclockwise rotation. The irradiated dose was set to 167 µSv with a measurement time of 100 s for all measurements. This reference dose was chosen as it is approximately in the middle of the low-dose range with regard to the dose variation [10]. Furthermore, 167 µSv has been used internally as a reference in the past for dosimetry measurements with Dosepix.
@@ IV-B Results for energy dependence
Fig. 2 shows the resulting response (measured by irradiated reference ) depending on the mean photon energy of the respective radiation qualities under an angle of incidence of . It should be noted that the conversion factors used for dose reconstruction have been optimized for Dpx 1 incorporated into the eye lens dosemeter prototype. As shown in [9], the energy dependent response differs slightly between individual Dosepix detectors and between read-out electronics. As this is not taken into account for the measurements presented in this work, the response is different compared to [9]. For future investigations, those effects should be considered by adjusting the dose conversion factors accordingly for each Dosepix. At low mean energies up to 48 keV (N-60), there are only minor deviations between the values measured at the different positions for both detectors. However, the differences in response increase as the mean photon energy increases, with the detector at the side measuring lower values than the detector on the eye. This is presumably caused by the scattering behavior of the setup, so that the discrepancy between the scattered radiation measured at the side of the Alderson phantom and on its eye increases with the mean photon energy of the radiation quality. Overall, there is a maximum difference of for Dpx 1 and for Dpx 2. These differences are small in comparison to the maximum energy dependent variation of between the values of N-250 and N-300 measured by Dpx 1 and between the values of N-40 and N-300 measured by Dpx 2. This proves that the intended measuring position of the active eye lens dosemeter on the side of the user’s head does not significantly impair the accuracy of the dosemeter.
@@ IV-C Results for angular dependence
Fig. 3 shows the results for the measurements at different angles of incidence. The response values measured with the detector on the eye show an increase for both positive and negative angles. This results from the increased scattering of the radiation in the Alderson phantom. This effect is less prominent for the measurements using N-30 as Dosepix behaves slightly differently at lower energies with respect to the angle of incidence [22]. On the side of the head, the resulting response values show a similar pattern, but the increase is higher for positive angles, because as the setup is turned in a positive direction, the detector on the side is closer to the beam source, i.e. closer to the X-ray tube, than the detector on the eye. Using the inverse squared relation between the expected dose and the distance to the radiation source, the difference in measured between the detector on the side and the detector on the eye is calculated for every angle. Fig. 3 shows the resulting values after correcting the results of Dpx 2 with respect to these positional differences. For irradiations with N-30 and N-60, the correction brings the values measured at the side (Dpx 2) closer to the values measured at the eye (Dpx 1). However, there are larger discrepancies for N-100 and N-200. As before, this can be explained by the scattering behavior of the phantom leading to an increase of scattered radiation with higher mean photon energies. Overall, the main cause for the differences between measuring at the side and measuring at the eye seems to be the distance to the beam source for lower mean photon energies, and scattering for higher mean photon energies.
The differences between the results measured on the side and on the eye without the correction regarding the distance are between and . This confirms the above conclusion that the measurement position on the side of the head, as intended for the active eye lens dosemeter, estimates the dose to the lens of the eye sufficiently well.
@@ V-A Lead Glass Pieces
The application of pieces of lead glass to the eye lens dosemeter prototype in order to mimic the effect of radiation safety glasses has been previously investigated in [23]. An image of the eye lens dosemeter prototype with the lead glass pieces is shown in Fig. 4. The size of the pieces was determined via intercept theorem by comparing the relation between the lead glass piece and Dosepix with the relation between the safety glasses and the eye lens. All values refer to the MAVIG BR126 protective glasses [24]. Such radiation safety glasses usually consist of frontal lenses made of lead glass as well as side protection. This is realized by two different pieces attached to the prototype. The frontal piece has a size of and is affixed to the front wall of the casing, shielding the Dosepix detector directly. The second piece is smaller () and is attached to the side of the dosemeter prototype facing away from the user’s head. Two different versions of the lead glass pieces were examined in terms of thickness. [23] has already examined the behavior of a frontal piece of thickness of 2.5 mm and a side piece of thickness 1.9 mm. However, in this work, a frontal piece with a thickness of 2.72 mm was used, as this reflects the MAVIG BR126 eye wear model more accurately. The thickness of the side piece remains at 1.9 mm.
@@ V-B Methods
To examine whether the lead glass pieces reproduce the shielding behavior of radiation safety glasses sufficiently well, several measurements were performed at the facilities of Physikalisch-Technische Bundesanstalt (PTB). A photograph of the setup is shown in Fig. 5. The eye lens dosemeter prototype with lead glass pieces was placed on the side of the head of an Alderson phantom [19] to account for the scattering behavior of a human head. For a comparison with the behind radiation safety glasses, two EYE-D dosemeters [25] were placed on each eye of the Alderson phantom behind the MAVIG BR126 protective glasses. EYE-D dosemeters are passive thermoluminescent dosemeters and are hereafter referred to as TLDs. For the analysis, only the outer TLDs were taken into account as they reflect the position of the eye lens better than the two inner ones. The inner TLDs provide substitutional data in case of a faulty measurement results from the outer TLDs. The eye lens dosemeter prototype was positioned in such a manner that its Dosepix detector was in the same plane as the TLDs.
The setup was irradiated with qualities of the N-series from N-15 to N-300 [21]. For N-30, N-60 and N-200, the irradiation angle was changed from 0∘ to , and . Negative angles indicate a clockwise rotation and positive angles a counterclockwise rotation. Therefore, the eye lens dosemeter prototype in this setup is further away from the beam source at negative angles and closer to the beam source at positive angles compared to the TLDs. The setup was rotated around a reference point according to ISO 29661 [20], which was set to the nasal root of the phantom, as indicated by the red laser cross in Fig. 5. A value for of 1 mSv was chosen for the radiation qualities N-15 to N-40 and 0.1 mSv for N-60 to N-300. These reference doses were chosen so that the response of the TLDs is well above their detection limit. The measurement time was 100 s for all measurements, except for N-15 with 200 s. Since the eye lens dosemeter prototype and the TLDs were shielded by lead glass, lower measured values than the reference dose were expected. A comparative measurement was performed without the radiation safety glasses and lead glass pieces with a reference dose of 1 mSv.
@@ V-D Results for energy dependence
A comparison of the normalized values measured by the eye lens dosemeter prototype and the TLDs depending on the mean photon energy of the irradiated photons is shown in Fig. 7. Fig. 7a shows the results of the measurements with lead glass pieces and radiation safety glasses and Fig. 7b the response values without the lead glass. The normalized values resulting from the measurements with lead glass pieces and radiation safety glasses (Fig. 7a) show an overall increase with the mean photon energy for both the eye lens dosemeter prototype and the TLDs. This is explained by the lead glass becoming more transmissive to radiation for increasing energies. For both detector types, there is a local maximum at 83 keV, i.e. for the irradiation with N-100. This is possibly explained by the K-edge of lead at 88 keV [26]. The mean photon energy of N-100 is just below this K-edge. The absorption at this energy is therefore less than at the next higher energetic irradiation quality N-120, which is with its mean energy of 100 keV just above the K-edge at 88 keV and therefore undergoes an increased absorption. Therefore, the measured dose at N-100 is higher than at N-120, possibly resulting in this local maximum.
Fig. 7a shows the response reconstructed directly from the measurement data of the eye lens dosemeter prototype as well as the corresponding corrected values. The latter were determined by applying certain correction factors to the response values in order to compensate for detector-related differences. Those factors were derived from the data without lead glass pieces or radiation safety glasses, as seen in Fig. 7b. Both detector types show a similar trend with respect to the mean photon energies. Nevertheless, it is evident that the response values of the eye lens dosemeter prototype fall mostly below the response values of the TLDs significantly. A possible explanation is given by the different energy dependencies of the two dosemeter types as well as their different positions on the Alderson phantom. To incorporate these differences in the results of the measurements with lead glass, correction factors for each energy are defined as the ratio of the values measured by the eye lens dosemeter prototype to the average of the left and right TLDs These factors are applied to the results of the active dosemeter prototype. This leads to the response values of the eye lens dosemeter prototype being closer to the values of the TLDs, as shown in Fig. 7a. It should be noted that specifying a reference dose behind the lead glass is difficult due to the unknown spectrum. Even after the correction, deviations up to between the results of the two detector types remain. There are several possible explanations for this. Since the surface of the lead glass pieces is smaller than that of the radiation safety glasses, the angular range in which the radiation is influenced by the lead glass is smaller. This leads to higher values measured by the TLDs which is consistent with the observations in Fig. 7a. Furthermore, the deposited energy spectrum of the irradiated quality changes when passing through lead glass as shown before in Fig. 6. The mean photon energy is therefore not identical when lead glass is present or not. This energy change results in small errors since the correction factors are energy dependent. However, this influence can be assumed to be negligibly small. Additionally, it is possible that the lead glass pieces might be of a slightly different thickness or material composition compared to the MAVIG BR126 glasses. Using the intercept theorem as a method to determine the size of the pieces might be insufficient. It is advisable to test the same methodology with other eye wear models and investigate whether similar behaviors occur.

### v2

# IEEE Copyright Notice

@@ Abstract
In this work, the effect of the measurement position on measurements of H(3) of a new active eye lens dosemeter prototype based on the Dosepix detector is examined. A comparison between measuring directly in front of the eye and measuring at the side of the head of an Alderson phantom showed no significant influence on the resulting H(3) for different radiation qualities and angles. In addition, to account for the absorption effect of radiation safety glasses, pieces of lead glass were attached to the front and side of the dosemeter. The challenging geometry of the human face, when additionally equipped with absorbing glasses and irradiated at different angles of radiation incidence, led to a complex irradiation setup, whereby the dosemeter cannot be placed at the real point of interest, i.e. directly in front of the eye. Corresponding effects and consequences for radiation protection measurements have been investigated by using a human like Alderson head phantom as well as thermoluminescent dosemeters (TLDs) and the active eye lens dosemeter prototype based on the Dosepix detector. For specific angles, the radiation bypassed the radiation safety glasses and lead glass pieces, leading to an increase in the measured H(3). Compared to TLDs behind radiation safety glasses, measurements with the lead glass shielded prototype resulted in lower values for H(3). Furthermore, the results did not reproduce previous findings where larger dose values were found for Dosepix behind lead glass pieces than for the TLDs behind radiation safety glasses. A possible reason might be that the dimensions of the lead glass pieces are not representative of the radiation safety glasses in front of the eye but, ultimately, it is not yet clear what the main reason for the deviation is. Therefore, it is advisable to test the same methodology in future investigations with other eyewear models and lead glass pieces to investigate whether similar behaviors occur.
@@
The relevant quantity in the context of eye lens dosimetry is the operational quantity . The International Commission on Radiation Units (ICRU) defines it as the equivalent personal dose at a depth of 3 mm in ICRU soft tissue [6, 7]. This quantity is used to estimate the dose deposited in the lens of the eye [8], i.e. the equivalent dose in the lens of the eye.
A prototype of an active dosemeter for the eye lens was developed at the Erlangen Centre for Astroparticle Physics (ECAP) as part of Friedrich-Alexander-Universität Erlangen-Nürnberg (FAU) [9]. This dosemeter is based on the hybrid pixel detector Dosepix and provides results within the IEC criteria in measurements [9]. However, it is unclear how the behavior of the dosemeter changes under more realistic conditions. So far, the eye lens dosemeter prototype has only been tested at the center front of a water cylinder phantom [9] in accordance with corresponding type test standards by IEC such as IEC 61526:2024 [10]. However, this does not represent the routine use since the dosemeter will be worn at the side of the head. This work examines the differences between those two measurement positions. Moreover, medical staff often use radiation safety glasses, which were not taken into account in previous investigations. In order to reproduce the shielding effect of such glasses, this work investigates the effectiveness of using lead glass pieces attached to the dosemeter.
@@ III The Eye Lens dosemeter
The prototype of the active eye lens dosemeter used for the measurements contains a Dosepix detector placed 2.5 mm behind the front wall of the 3D-printed casing. To ensure a stable position and prevent possible damage to the sensor, a PMMA ring with a diameter of 18 mm and a thickness of 2.5 mm is fixed around the detector. The read-out hardware is connected to the Dosepix detector via a flat cable. The eye lens dosemeter is intended to be worn at the side of the head, so that the Dosepix detector faces to the front. Detailed information as well as the results of performance tests are provided in [9].
During the measurement process, each detector pixel compares incoming energy depositions to 16 freely adjustable energy bins and sorts them accordingly. The lowest bin edge is set to 12 keV. The highest bin is used as an overflow bin and contains all events for which the deposited energy exceeds 150 keV. The counts in each bin are summed up to the total number of entries , which is then multiplied with a corresponding dose conversion factor . The weighted sum over all 16 bins determines the equivalent dose of the eye lens .
The dose conversion factors used to obtain the dose from the detector signal were determined by a combination of measurements and radiation transport simulations and yield reliable results for photon radiation fields with energies up to about 250 keV [9]. The determination of these dose conversion factors represent the calibration of the device. The statistical uncertainty of the dose is estimated from Poisson statistics with the law of uncertainty propagation under the assumption of independent [15, 16].
In the following, only statistical uncertainties are considered. Other uncertainties, e.g. resulting from the positioning, have no significant influence on the results due to the detailed adjustment at the measurement facilities.
It should be noted that there is only a single set of dose conversion factors for all Dosepix detectors, which was derived from measurement data from a specific detector (Dpx 1 in the following chapters) [9]. Since there are variations between individual detectors, this leads also to different results regarding reconstructed dose. However, these differences are within the limits of 0.71 and 1.67 stated in IEC 61526:2024 [10], as shown in [9]. These limits are not symmetrical around the ideal response of 1.0, since an underestimation of the deposited dose has more severe consequences than an overestimation, which is taken into account for the definition of the response limits.
@@ IV-A Methods
Fig. 1 shows an image of the setup for the measurements at the facilities of Physikalisch-Technische Bundesanstalt (PTB). To compare the values of measured on the eye and at the side of the head, two Dosepix detectors (numbered as Dpx 1 and Dpx 2) were placed on an Alderson phantom [19]. This phantom allows for a more representative position of the dosemeter as it would be worn by the user. It should be noted that neither of the two detectors is installed in a dosemeter prototype, but is directly connected to the read-out hardware. To minimize the amount of material in front of the phantom, a connector extension is installed between Dosepix and its read-out hardware. Both detectors are in the same plane perpendicular to the beam axis. The reference point according to ISO 29661 [20] is set to the nasal root of the Alderson phantom at 2.5 m distance to the X-ray tube. The detector at the side is placed at a distance of 7.0 cm to the reference point, the detector on the eye at a distance of 2.5 cm. The detector on the eye was placed as close to the surface of the phantom as possible. Due to the hardware, a distance of 1.5 cm was between sensor and surface.
The setup was irradiated with radiation qualities of the narrow-spectrum series (N-series) from N-20 to N-300 (with mean energies from 16 keV to 248 keV), with one irradiation per quality. Detailed information regarding the qualities’ characteristics is given in ISO 4037-1 [21]. This series of radiation qualities was chosen for comparability reasons with previous measurements [9, 16]. The whole setup was illuminated homogeneously. To account for possible differences between the two detectors, the measurements were repeated where the detectors, including the respective hardware, were interchanged. Therefore, both detectors measured at both positions. Additionally, the angle of irradiation was varied between 0∘ to , and for the measurements with N-30, N-60, N-100 and N-200. Negative angles denote a clockwise rotation and positive angles a counterclockwise rotation. It should be noted that only the rotation regarding left/right anterior oblique view was examined since the setup does not enable a rotation in cranial/caudal direction. Since the latter case is common in the medical field, future studies should examine the behavior for angles in cranial/caudal directions. However, the influence is expected to be small since the angular dependency of the Dosepix detector is identical for horizontal and vertical rotation. The irradiated dose was set to 167 µSv with a measurement time of 100 s for all measurements. This reference dose was chosen as it is approximately in the middle of the low-dose range with regard to the dose variation [10]. Furthermore, 167 µSv has been used internally as a reference in the past for dosimetry measurements with Dosepix.
@@ IV-B Results for energy dependence
Fig. 2 shows the resulting response (measured by irradiated reference ) depending on the mean photon energy of the respective radiation qualities under an angle of incidence of .
As shown in [9], the energy dependent response differs slightly between individual Dosepix detectors and between read-out electronics. As this is not taken into account for the measurements presented in this work, the response is different compared to [9]. For future investigations, those effects should be considered by adjusting the dose conversion factors accordingly for each Dosepix using a certain method that will be established in future research.
At low mean energies up to 48 keV (N-60), there are only minor deviations between the values measured at the different positions for both detectors. However, the differences in response increase as the mean photon energy increases, with the detector at the side measuring lower values than the detector on the eye. This is presumably caused by the scattering behavior of the setup, so that the discrepancy between the scattered radiation measured at the side of the Alderson phantom and on its eye increases with the mean photon energy of the radiation quality. Overall, there is a maximum difference of for Dpx 1 and for Dpx 2. These differences are small in comparison to the maximum energy dependent variation of between the values of N-250 and N-300 measured by Dpx 1 and between the values of N-40 and N-300 measured by Dpx 2. Given the stated differences, this suggests that the intended measuring position of the active eye lens dosemeter on the side of the user’s head does not significantly impair the accuracy of the dosemeter.
@@ IV-C Results for angular dependence
Fig. 3 shows the results for the measurements at different angles of incidence. The response values measured with the detector on the eye show an increase compared to the value measured at 0∘ for both positive and negative angles. Since the detectors are not shadowed by the phantom even at 75∘, this results from the increased scattering of the radiation in the Alderson phantom. This effect is less prominent for the measurements using N-30 as Dosepix behaves slightly differently at lower energies with respect to the angle of incidence [22]. On the side of the head, the resulting response values show a similar pattern, but the increase is higher for positive angles, because as the setup is turned in a positive direction, the detector on the side is closer to the beam source, i.e. closer to the X-ray tube, than the detector on the eye. Using the inverse squared relation between the expected dose and the distance to the radiation source, the difference in measured between the detector on the side and the detector on the eye is calculated for every angle. Fig. 3 shows the resulting values after correcting the results of Dpx 2 with respect to these positional differences. For irradiations with N-30 and N-60, the correction brings the values measured at the side (Dpx 2) closer to the values measured at the eye (Dpx 1). However, there are larger discrepancies for N-100 and N-200. As before, this can be explained by the scattering behavior of the phantom leading to an increase of scattered radiation with higher mean photon energies. Overall, the main cause for the differences between measuring at the side and measuring at the eye seems to be the distance to the beam source for lower mean photon energies, and scattering for higher mean photon energies.
The differences between the results measured on the side and on the eye without the correction regarding the distance are between (N-100 at -60∘) and (N-30 at 60∘). This confirms the above conclusion that the measurement position on the side of the head, as intended for the active eye lens dosemeter, estimates the dose to the lens of the eye sufficiently well.
@@ V-A Lead Glass Pieces
The application of pieces of lead glass to the eye lens dosemeter prototype in order to mimic the effect of radiation safety glasses has been previously investigated in [23]. An image of the eye lens dosemeter prototype with the lead glass pieces is shown in Fig. 4. The size of the pieces was determined via intercept theorem [24] by comparing the relation between the lead glass piece and Dosepix with the relation between the safety glasses and the eye lens. All values refer to the MAVIG BR126 protective glasses [25]. Such radiation safety glasses usually consist of frontal lenses made of lead glass as well as side protection. This is realized by two different pieces attached to the prototype. The frontal piece has a size of and is affixed to the front wall of the casing, shielding the Dosepix detector directly. The second piece is smaller () and is attached to the side of the dosemeter prototype facing away from the user’s head. Two different versions of the lead glass pieces were examined in terms of thickness. The behavior of a frontal piece of thickness of 2.5 mm and a side piece of thickness 1.9 mm was already examined in [23]. However, in this work, a frontal piece with a thickness of 2.72 mm was used, as this reflects the MAVIG BR126 eye wear model more accurately. The thickness of the side piece remains at 1.9 mm.
@@ V-B Methods
To examine whether the lead glass pieces reproduce the shielding behavior of radiation safety glasses sufficiently well, several measurements were performed at the facilities of Physikalisch-Technische Bundesanstalt (PTB). A photograph of the setup is shown in Fig. 5. The eye lens dosemeter prototype (containing Dpx 1) with lead glass pieces was placed on the side of the head of an Alderson phantom [19] to account for the scattering behavior of a human head. For a comparison with the behind radiation safety glasses, two EYE-D dosemeters [26] were placed on each eye of the Alderson phantom behind the MAVIG BR126 protective glasses. EYE-D dosemeters are passive thermoluminescent dosemeters and are hereafter referred to as TLDs. For the analysis, only the outer TLDs were taken into account as they reflect the position of the eye lens better than the two inner ones. The inner TLDs provide substitutional data in case of a faulty measurement results from the outer TLDs. The eye lens dosemeter prototype was positioned in such a manner that its Dosepix detector was in the same plane as the TLDs.
The setup was irradiated with qualities of the N-series from N-15 to N-300 [21]. For N-30, N-60 and N-200, the irradiation angle was changed from 0∘ to , and . Negative angles indicate a clockwise rotation and positive angles a counterclockwise rotation. Therefore, the eye lens dosemeter prototype in this setup is further away from the beam source at negative angles and closer to the beam source at positive angles compared to the TLDs. The differences in distances were not corrected using the inverse squared relation as before, since this corresponds to a more realistic comparison between the dose measured at the eye and at the eye lens dosemeter prototype. The setup was rotated around a reference point according to ISO 29661 [20], which was set to the nasal root of the phantom, as indicated by the red laser cross in Fig. 5. A value for of 1 mSv was chosen for the radiation qualities N-15 to N-40 and 0.1 mSv for N-60 to N-300. These reference doses were chosen so that the response of the TLDs is well above their detection limit. The measurement time was 100 s for all measurements, except for N-15 with 200 s. Since the eye lens dosemeter prototype and the TLDs were shielded by lead glass, lower measured values than the reference dose were expected. A comparative measurement was performed without the radiation safety glasses and lead glass pieces with a reference dose of 1 mSv.
@@ V-D Results for energy dependence
A comparison of the normalized values measured by the eye lens dosemeter prototype and the TLDs depending on the mean photon energy of the irradiated photons is shown in Fig. 7. Fig. 7a shows the results of the measurements with lead glass pieces and radiation safety glasses and Fig. 7b the response values without the lead glass. All irradiations were performed at an angle of 0∘.
To compare the eye lens dosemeter prototype and the TLDs, measurements were performed without any lead glass shielding (see Fig. 7b). Both detector types show a similar trend with respect to the mean photon energies. Nevertheless, it is evident that the response values of the eye lens dosemeter prototype mostly fall significantly below the response values of the TLDs. A possible explanation is given by the different energy dependencies of the two dosemeter types as well as their different positions on the Alderson phantom. To incorporate these differences in the results of the measurements with lead glass, correction factors for each energy are defined as the ratio of the values measured by the eye lens dosemeter prototype to the average of the left and right TLDs.
Fig. 7a shows the normalized values resulting from the measurements with lead glass pieces and radiation safety glasses, with the response reconstructed directly from the measurement data of the eye lens dosemeter prototype as well as the corresponding corrected values. The latter were determined by applying correction factors calculated from the data measured without lead glass shielding, as explained above. This compensates for detector-related differences as well as different measurement positions of TLDs and eye lens dosemeter prototype. This allows the focus to be put on the influence of the lead glass.
The normalized values resulting from the measurements with lead glass pieces and radiation safety glasses in Fig. 7a show an overall increase with the mean photon energy for both the eye lens dosemeter prototype and the TLDs. This is explained by the lead glass becoming more transmissive to radiation for increasing energies. For both detector types, there is a local maximum at 83 keV, i.e. for the irradiation with N-100. This is possibly explained by the K-edge of lead at 88 keV [27]. The mean photon energy of N-100 is just below this K-edge. The absorption at this energy is therefore less than at the next higher energetic irradiation quality N-120, which is with its mean energy of 100 keV just above the K-edge at 88 keV and therefore undergoes an increased absorption. Therefore, the measured dose at N-100 is higher than at N-120, possibly resulting in this local maximum.
After applying the correction factors, the response values of the eye lens dosemeter prototype are closer to the values of the TLDs, as shown in Fig. 7a.
Even after the correction, deviations up to between the results of the two detector types remain. There are several possible explanations for this. Since the surface of the lead glass pieces is smaller than that of the radiation safety glasses, the angular range in which the radiation is influenced by the lead glass is smaller. This leads to higher values measured by the TLDs which is consistent with the observations in Fig. 7a.
Furthermore, the deposited energy spectrum of the irradiated quality changes when passing through lead glass as shown before in Fig. 6. The mean photon energy is therefore not identical when lead glass is present or not. This energy change results in small errors since the correction factors are energy dependent. However, this influence can be assumed to be negligibly small.
Additionally, it is possible that the lead glass pieces might be of a slightly different thickness or material composition compared to the MAVIG BR126 glasses. Using the intercept theorem as a method to determine the size of the pieces might be insufficient. It is advisable to test the same methodology with other eye wear models and investigate whether similar behaviors occur.